# EP241021a: A catastrophic collapse/merger of compact star binary leading to the formation of a remnant millisecond magnetar?

### v1

@@ 1 Introduction
The Einstein Probe (EP; also known as Tian-Guan in China) satellite is a mission designed to detect transient phenomena in the soft X-ray sky. Two telescopes onboard include the Wide-field X-ray Telescope (WXT), monitoring the sky with its large 3600 deg2 field of view, and the Follow-up X-ray Telescope (FXT), which can quickly characterize and precisely localize the transients triggered onboard or commanded uplink (Yuan et al., 2015, 2022, 2025). Since its launch on 9 January 2024, EP has successfully captured more than 130 new fast X-ray transients (FXRTs), around 20 of which were found to be associated with a gamma-ray burst (GRB), indicating a potential link between the FXRTs and GRBs. However, for many FXRTs, actually no GRB counterpart is detected, even though they are located in the field of view of gamma-ray telescopes that are sensitive enough to observe a typical GRB. Even in the association case such as EP240315a/GRB 240315C, the durations and temporal behaviors of the associated FXRT and GRB could still be very different from each other (Liu et al., 2025; Levan et al., 2024). This indicates that the FXRT emission may not simply be considered to be the extension of GRB emission into the low-luminosity range or into the low-energy band.
@@
It is also striking that some FXRTs can further lead to the discovery of an unusual optical transient, e.g., as first reported, EP240414a was found to be followed by an unusual broad-lined Type Ic supernova (SN Ic-BL): SN 2024gsa (Srivastav et al., 2025; van Dalen et al., 2025; Sun et al., 2024). This situation is similar to that appearing in the XRF 060218/SN 2006aj association event. According to its redshift of $z=0.401$, the luminosity of SN 2024gsa can be inferred to be as high as $\sim 2\times 10^{43}\rm erg~{}s^{-1}$, which requires a mass of $0.74^{+0.05}_{-0.04}M_{\odot}$ of 56Ni if it is powered by radioactivity as usual. This mass of 56Ni is actually unusually high even for the population of GRB-associated SNe (e.g., Woosley & Bloom, 2006; Hjorth & Bloom, 2012; Cano, 2013; Prentice et al., 2016; Lü et al., 2018). In addition, before SN 2024gsa, another significant optical light curve bump was found to appear two days after the FXRT, which rises very quickly to a peak absolute magnitude of $-21$ AB mag within a few days. The origin of this early bump is debated (Sun et al., 2024; Hamidani et al., 2025). Very recently, such a double-bump optical light curve was also discovered following FXRT EP250108a (Rastinejad et al., 2025; Eyles-Ferris et al., 2025; Srinivasaragavan et al., 2025; Li et al., 2025), where a more luminous SN (e.g., SN 2025kg) was accompanied by weaker X-ray emission.
@@
The SNe following FXRTs EP240414a and EP250108a are both too luminous to be accounted for solely by the traditional radioactive power, which makes them potentially connected to GRBs as well as superluminous supernovae (SLSNe). Therefore, a central engine could be expected to play a crucial role in driving these FXRTs, since such an engine (in the particular sense, a rapidly rotating magnetar) was usually used to explain the phenomena of GRBs and SLSNe (e.g., Dai & Lu, 1998; Zhang & Mészáros, 2001; Yu & Dai, 2007; Kasen & Bildsten, 2010; Metzger et al., 2015; Yu et al., 2017; Margalit et al., 2018; Fiore et al., 2025). Furthermore, the big bump on a timescale of several days, which is before the SN peak, could provide more detailed clues to the explosion processes and environments of the FXRTs. First, the bump could be the result of the cooling of SN ejecta and dense circum-stellar material that were previously heated by the SN shock (Sun et al., 2024; van Dalen et al., 2025; Srinivasaragavan et al., 2025), in particular, when the SN shock can be energized by the powerful wind of a magnetar (Zhang et al., 2022). Nevertheless, it was also suggested that the bump could have a non-thermal origin, e.g., due to a shock refreshed by itself (Srivastav et al., 2025) or a shock driven by an off-axis cocoon (Hamidani et al., 2025; Zheng et al., 2025).
@@
Very recently, a significant light curve bump has been further discovered from EP241021a, which appeared ten days after the FXRT emission (Busmann et al., 2025; Gianfagna et al., 2025; Shu et al., 2025; Yadav et al., 2025). The timescale of this bump is just shorter than the typical timescale of SNe of about several tens of days, but relatively longer than the cooling time of SN ejecta, which is on the order of a few days. Meanwhile, the absolute peak magnitude of the optical bump can be set at $M_{\mathrm{r}}\approx-22$ mag, according to its relatively high redshift $z=0.748$ (Pugliese et al., 2024; Perez-Fournon et al., 2024; Zheng et al., 2024). This luminosity and timescale of the optical bump make it resemble some mysterious fast blue optical transients (FBOTs; Drout et al., 2014) such as AT 2018cow, although the optical color of EP 241021a is redder. It is worth mentioning that AT 2018cow was prominent in the FBOT population due to its abundant X-ray emission (Perley et al., 2019; Rivera Sandoval et al., 2018; Margutti et al., 2019), while the optical bump of EP241021a is also accompanied by an X-ray rebrightening that is completely different from EP240414a and EP250108a. The evolution behavior of the X-ray emission of AT 2018cow (Margutti et al., 2019) and the possible quasi-periodic oscillation in it (Pasham et al., 2022) further shows a strong potential connection between this X-ray emission and the spin down of a millisecond magnetar. Therefore, in this Letter, we focus on the possibility that the multi-wavelength bump of EP241021a was contributed by an explosion-type transient, in view that some other types of model have been previously proposed including a refreshed shock (Busmann et al., 2025), a structured jet as well as its cocoon (Gianfagna et al., 2025), and an intermittent energized shock due to repeating partial tidal disruption event involving an intermediate-mass black hole (Shu et al., 2025).
@@ 2 Observation of EP241021a
EP241021a was initially detected by the WXT onboard EP at 05:07:56 UTC with a time-averaged X-ray flux of $3.31^{+1.26}_{-0.86}\times 10^{-10}\mathrm{erg~{}cm^{-2}~{}s^{-1}}$ in the 0.5–4.0 keV. The transient has a duration of $\sim 100~{}\mathrm{s}$ and a relatively hard X-ray spectrum. No high-energy gamma-ray counterpart to EP241021a was detected by Fermi-GBM or Konus-Wind, despite targeted and blind searches around the trigger time. After 36.5 hours of the initial detection, EP241021a was observed by the FXT on board EP. A total of 12 follow-up observations were conducted while XMM-Newton gives an upper limit estimated at $T_{0}+89$ days. As shown in Figure 1, an X-ray light curve was revealed, which contains an initial decay, a plateau with a slight rise between $T_{0}+3$ and $T_{0}+10$ days, and a second decay roughly following a behavior of $\sim t^{-2}$ until an abrupt decline around $T_{0}+80$ days.
@@
Optical photometry of EP241021a was first reported by Nordic Optical Telescope at R.A. = 01:55:23.41, Dec. = +05:56:18.01 with an uncertainty of $0.5^{\prime\prime}$ (Fu et al., 2024) and was confirmed by other earlier observations. The following observations were performed using multiple instruments (e.g., Thai Robotic Telescope, Katzman Automatic Imaging Telescope, Liverpool Telescope, ALT-50A&-100C, HMT-0.5m Telescope, Gran Telescopio Canarias, Fraunhofer Telescope Wendelstein, Hobby-Eberly Telescope, Very Large Telescope, The Swift Ultra-Violet Optical Telescope, Calar Alto Telescope, Large Binocular Telescope; Busmann et al., 2025; Gianfagna et al., 2025; Shu et al., 2025). As presented in Figure 2, the initial decay of the optical light curves could basically trace the early X-ray emission, while a significant bump appears in all bands coinciding with the X-ray plateau/rebrightening. Besides the first bump around $T_{0}+10$ days, a second bump seems to exist at $T_{0}+50$ days although the signal is still ambiguous. The optical spectroscopy of EP241021a was conducted by Gran Telescopio Canarias (Pérez-Fournon et al., 2024), Keck I 10 m telescope (Zheng et al., 2024), Very Large Telescope (Pugliese et al., 2024), and Hobby-Eberly Telescope (Busmann et al., 2025). The detected narrow emission lines of [O II] and [O III] are likely from the host galaxy and give the redshift of $z=0.748$, while no SN-like features are identified in these spectra.
@@
Finally, radio observations were carried out with VLA, MeerKAT, VLBA, e-MERLIN, uGMRT, and ATCA simultaneously. The temporal evolution of the radio emission can be roughly described by a power-law rise and decline with a peak shifting toward lower frequencies over time, as displayed in Figure 3.
@@
Figure 2: Observed optical light curves of EP241021a (solid circles) and the model fittings (lines). The data are taken from Busmann et al. (2025) and Shu et al. (2025) but with an extra correction of $E(B-V)=0.045~{}\mathrm{mag}$ and $R_{{V}}=0.31$ corresponding to the Galactic extinction (Schlafly & Finkbeiner, 2011).
Figure 3: Radio emission of EP241021a and the fitting with a jet shock emission. The data are taken from Shu et al. (2025).
@@ 3.1 An explosion transient emission
The peak bolometric luminosity of the optical bump at 10 days can be calculated to $L_{\rm p}\sim 10^{44}\mathrm{erg~{}s^{-1}}$ with a black body assumption. The blackbody temperature can be estimated to be within the range of $T=(1+z)h\nu/(2.82k_{\mathrm{B}})=(0.7\sim 1.7)\times 10^{4}~{}\mathrm{K}$, where the spectral peak frequency is considered to be located in the observational range of $(4\sim 10)\times 10^{14}~{}\mathrm{Hz}$, $k_{\mathrm{B}}$ and $h$ are the Boltzmann constant and the Planck constant, respectively. Then, the radius of the emitting photosphere can be estimated as
@@
$\simeq 3.8\times 10^{15}\left(\frac{L_{\rm p}}{10^{44}~{}\mathrm{erg~{}s^{-1}}}\right)^{1/2}\left(\frac{T}{10^{4}~{}\mathrm{K}}\right)^{-2}\mathrm{cm},$
@@
$\simeq 6.5\left(\frac{\kappa}{10~{}\mathrm{{cm}^{2}g^{-1}}}\right)^{1/2}\left(\frac{M_{\mathrm{ej}}}{0.02~{}M_{\odot}}\right)^{1/2}\left(\frac{v_{\mathrm{ej}}}{0.25~{}c}\right)^{-1/2}~{}\mathrm{day},$
@@
where the initial value of the spin-down luminosity and the spin-down timescale are expressed as usual as $L_{\mathrm{sd,i}}=10^{47}B_{\mathrm{p,14}}^{2}P_{\mathrm{i,-3}}^{-4}~{}\mathrm{erg~{}s^{-1}}$ and $t_{\mathrm{sd}}=2\times 10^{5}~{}(1+z)B_{\mathrm{p,14}}^{-2}P_{\mathrm{i,-3}}^{2}~{}\mathrm{s}$ with $B_{\mathrm{p}}$ and $P_{\mathrm{i}}$ being the magnetic field strength and initial spin period of the magnetar, respectively. The convention $Q_{x}=Q/10^{x}$ is adopted in cgs units. The coefficient $\xi=1-\exp(-\tau_{\mathrm{x}})$ is introduced to represent the fraction of the spin-down energy absorbed by the explosion ejecta, where $\tau_{\mathrm{x}}=3\kappa_{\mathrm{x}}M_{\mathrm{ej}}/(4\pi v_{\mathrm{ej}}^{2}t^{2})$ is the optical depth for X-rays with $\kappa_{\mathrm{x}}$ being the effective X-ray opacity. It is considered that the spin-down energy released from the magnetar is initially in the form of a Poynting flux and subsequently carried out by an ultra-relativistic electron-positron wind. This energy conversion occurs mainly through magnetic reconnection processes, which can simultaneously produce X-ray emission. When the magnetar wind catches up and collides with the previous ejecta, a reverse shock can be formed in the wind to convert its kinetic energy into heat. The hot electron-positron plasma in the reverse-shocked wind (i.e., the magnetar wind nebula; MWN) can release its internal energy through the synchrotron and inverse-Compton processes of relativistic leptons. As a result, the emitted photons would be mainly in X-rays (Metzger et al., 2014; Yu et al., 2019) and, therefore, the X-ray opacity is involved in the absorption coefficient.
@@
[…]
@@
Here, it should be pointed out that, as the ejecta expands, the X-ray emission generated from the MWN can increasingly penetrate the ejecta to be detected. Thus, a rebrightening of the X-ray emission accompanying the optical bump can be naturally understood. Such a feature makes EP241021a intrinsically different from the other analogous transient events such as EP240414a and EP250108a. In other words, the modeling of the X-ray emission of EP241021a would play a crucial role in identifying its origin.

### v2

@@ 1 Introduction
The Einstein Probe (EP; also known as Tian-Guan in China) satellite is a mission designed to detect transient phenomena in the soft X-ray sky. Two telescopes onboard include the Wide-field X-ray Telescope (WXT), monitoring the sky with its large 3600 deg2 field of view, and the Follow-up X-ray Telescope (FXT), which can quickly characterize and precisely localize the transients triggered onboard or commanded uplink (Yuan et al., 2015, 2022, 2025). Since its launch on 9 January 2024, EP has successfully captured more than 130 new fast X-ray transients (FXRTs), around 20 of which were found to be associated with a gamma-ray burst (GRB), indicating a potential link between the FXRTs and GRBs. However, for many FXRTs, actually no GRB counterpart is detected, even though they are located in the field of view of gamma-ray telescopes that are sensitive enough to observe a typical GRB (Sun et al., 2024; Busmann et al., 2025; Ravasio et al., 2025; Zhang et al., 2025). Even in the association case such as EP240315a/GRB 240315C, the durations and temporal behaviors of the associated FXRT and GRB could still be very different from each other (Liu et al., 2025; Levan et al., 2024). This indicates that the FXRT emission may not simply be considered to be the extension of GRB emission into the low-luminosity range or into the low-energy band.
@@
It is also striking that some FXRTs can further lead to the discovery of an unusual optical transient, e.g., as first reported, EP240414a was found to be followed by an unusual broad-lined Type Ic supernova (SN Ic-BL): SN 2024gsa (Srivastav et al., 2025; van Dalen et al., 2025; Sun et al., 2024). This situation is similar to that appearing in the XRF 060218/SN 2006aj association event. According to its redshift of $z=0.401$, the luminosity of SN 2024gsa can be inferred to be as high as $\sim 2\times 10^{43}\rm erg~s^{-1}$, which requires a mass of $0.74^{+0.05}_{-0.04}M_{\odot}$ of 56Ni if it is powered by radioactivity as usual. This mass of 56Ni lies at the upper edge of the range of $M_{\rm Ni}$ of GRB-associated SNe (i.e., $\sim 0.2-0.6M_{\odot}$ [^1] The measurements of nickle masses for GRB-associated SNe are usually degenerate with the modeling of the simultaneous jet afterglow emission, which can lead to a systematic uncertainty of the nickle masses.; Cano, 2013; Prentice et al., 2016; Cano et al., 2017; Lü et al., 2018), indicating that SN 2025gsa requires an unusually large amount of 56Ni synthesis. It should be noted that the range of $\sim 0.2-0.6M_{\odot}$ is already several times higher than that of normal core-collapse SNe which is usually no more than $\sim 0.2M_{\odot}$ (Woosley & Bloom, 2006; Hjorth & Bloom, 2012). In addition, before SN 2024gsa, another significant optical light curve bump was found to appear two days after the FXRT, which rises very quickly to a peak absolute magnitude of 21-21 AB mag within a few days. The origin of this early bump is debated (Sun et al., 2024; Hamidani et al., 2025). Very recently, such a double-bump optical light curve was also discovered following FXRT EP250108a (Rastinejad et al., 2025; Eyles-Ferris et al., 2025; Srinivasaragavan et al., 2025; Li et al., 2025), where a more luminous SN (i.e., SN 2025kg) was accompanied by weaker X-ray emission.
@@
The SNe following FXRTs EP240414a and EP250108a are both too luminous to be accounted for solely by the traditional radioactive power, which makes them potentially connected to GRBs as well as superluminous supernovae (SLSNe). Therefore, a central engine could be expected to play a crucial role in driving these FXRTs, since such an engine (in the particular sense, a rapidly rotating magnetar) was usually used to explain the phenomena of GRBs and SLSNe (e.g., Dai & Lu, 1998; Zhang & Mészáros, 2001; Yu & Dai, 2007; Kasen & Bildsten, 2010; Metzger et al., 2015; Yu et al., 2017; Margalit et al., 2018; Fiore et al., 2025). Furthermore, the big bump on a timescale of several days, which is before the SN peak, could provide more detailed clues to the explosion processes and environments of the FXRTs. First, the bump could be the result of the cooling of SN ejecta and dense circum-stellar material that were previously heated by the SN shock (Sun et al., 2024; van Dalen et al., 2025; Srinivasaragavan et al., 2025), in particular, when the SN shock can be energized by the powerful wind of a magnetar (Zhang et al., 2022). Nevertheless, it was also suggested that the bump could have a non-thermal origin, e.g., due to a shock driven by a jet and refreshed by the jet tail or due to a shock driven by an off-axis cocoon (Hamidani et al., 2025; Zheng et al., 2025). In the former scenario, it is assumed that the velocity stratification of the jet enables slower shells to catch up with the decelerating shock front, thereby re-energizing the front and inducing a delayed brightening (Rees & Mészáros, 1998; Sari & Mészáros, 2000; Srivastav et al., 2025).
@@
Very recently, a significant light curve bump has been further discovered from EP241021a, which appeared ten days after the FXRT emission (Busmann et al., 2025; Gianfagna et al., 2025; Shu et al., 2025; Yadav et al., 2025). The timescale of this bump is just shorter than the typical timescale of SNe of about several tens of days, but relatively longer than the cooling time of SN ejecta, which is on the order of a few days. Meanwhile, the absolute peak magnitude of the optical bump can be set at $M_{\mathrm{r}}\approx-22$ mag, according to its relatively high redshift $z=0.748$ (Pugliese et al., 2024; Perez-Fournon et al., 2024; Zheng et al., 2024). This luminosity and timescale of the optical bump make it resemble some mysterious fast blue optical transients (FBOTs; Drout et al., 2014) such as AT 2018cow, although the optical color of EP 241021a is redder. It is worth mentioning that AT 2018cow was prominent in the FBOT population due to its abundant X-ray emission (Perley et al., 2019; Rivera Sandoval et al., 2018; Margutti et al., 2019), while the optical bump of EP241021a is also accompanied by an X-ray rebrightening that is completely different from EP240414a and EP250108a. The evolution behavior of the X-ray emission of AT 2018cow (Margutti et al., 2019) and the possible quasi-periodic oscillation in it (Pasham et al., 2022) further shows a strong potential connection between this X-ray emission and the spin down of a millisecond magnetar. Therefore, in this paper, we focus on the possibility that the multi-wavelength bump of EP241021a was contributed by an explosion-type transient, in view that some other types of model have been previously proposed including a refreshed shock (Busmann et al., 2025), a structured jet as well as its cocoon (Gianfagna et al., 2025), and an intermittent energized shock due to repeating partial tidal disruption event involving an intermediate-mass black hole (Shu et al., 2025).
@@ 2 Observation of EP241021a
EP241021a was initially detected by the WXT onboard EP at 05:07:56 UTC with a time-averaged X-ray flux of $3.31^{+1.26}_{-0.86}\times 10^{-10}\mathrm{erg~cm^{-2}~s^{-1}}$ in the 0.5–4.0 keV photon energy range. The transient has a duration of $\sim 90~\mathrm{s}$ and a relatively hard X-ray spectrum with a photon index of $1.8^{+0.57}_{-0.54}$ (Shu et al., 2025). No high-energy gamma-ray counterpart to EP241021a was detected by Fermi-GBM (Burns et al., 2024) or Konus-Wind (Svinkin et al., 2024), despite both targeted and blind searches around the trigger time. Specifically, Fermi-GBM places a $3\sigma$ upper limit on the fluence in the $30-400~\mathrm{keV}$ band of $9.8\times 10^{-7}\mathrm{erg~cm^{-2}}$ (Gianfagna et al., 2025) and Konus-Wind reports a $90\%$ confidence upper limit on the $20-1,500~\mathrm{keV}$ peak flux of $2.5\times 10^{-7}\mathrm{erg~cm^{-2}s^{-1}}$ in a timescale of $2.994~\mathrm{s}$ (Svinkin et al., 2024). These gamma-ray non-detections constrain the spectral peak energy of EP 241021a to $E_{\mathrm{p}}\lesssim 100~\mathrm{keV}$ and the isotropic-equivalent energy to $E_{\mathrm{iso}}\lesssim 2\times 10^{51}~\mathrm{erg}$, where a Band function extrapolated from the WXT detection is assumed. After 36.5 hours of the initial detection, EP241021a was observed by the FXT on board EP, with a total of 12 observations conducted (Shu et al., 2025). Follow-up observations were also performed with XMM-Newton, yielding an upper limit at $T_{0}+89$ days (Shu et al., 2025). As shown in Figure 1, an X-ray light curve was revealed, which contains a hint of an initial decay, a plateau with a possible slight rise between $T_{0}+3$ and $T_{0}+10$ days, and a second decay roughly following a behavior of $\sim t^{-2}$ until an abrupt decline around $T_{0}+80$ days.
@@
Optical photometry of EP241021a was first reported by Nordic Optical Telescope at R.A. = 01:55:23.41, Dec. = +05:56:18.01 with an uncertainty of $0.5^{\prime\prime}$ (Fu et al., 2024) and was confirmed by other earlier observations. The following observations were performed using multiple instruments (e.g., Thai Robotic Telescope, Katzman Automatic Imaging Telescope, Liverpool Telescope, ALT-50A&-100C, HMT-0.5m Telescope, Gran Telescopio Canarias, Fraunhofer Telescope Wendelstein, Hobby-Eberly Telescope, Very Large Telescope, The Swift Ultra-Violet Optical Telescope, Calar Alto Telescope, Large Binocular Telescope; Busmann et al., 2025; Gianfagna et al., 2025; Shu et al., 2025). As presented in Figure 2, the initial decay of the optical light curves could basically trace the early X-ray emission, while a significant bump appears in all bands coinciding with the X-ray plateau/rebrightening. Besides the first bump around $T_{0}+10$ days, a second bump seems to exist at $T_{0}+50$ days (see the argument in Shu et al. 2025) although the signal is still ambiguous. The optical spectroscopy of EP241021a was conducted by Gran Telescopio Canarias (Pérez-Fournon et al., 2024), Keck I 10 m telescope (Zheng et al., 2024), Very Large Telescope (Pugliese et al., 2024), and Hobby-Eberly Telescope (Busmann et al., 2025). The detected narrow emission lines of [O II] and [O III] are likely from the host galaxy and give the redshift of $z=0.748$, while no SN-like features are identified in these spectra.
@@
Finally, radio observations were carried out with ATCA (Gianfagna et al., 2025; Yadav et al., 2025), e-MERLIN (Gianfagna et al., 2025; Yadav et al., 2025), uGMRT, ALMA (Gianfagna et al., 2025), VLA, MeerKAT, and VLBA (Shu et al., 2025), simultaneously. The temporal evolution of the radio emission can be roughly described by a power-law rise and decline with a peak shifting toward lower frequencies over time, as displayed in Figure 3.
@@
Figure 2: Observed optical light curves of EP241021a (solid circles) and the model fittings (lines). The data are taken from Busmann et al. (2025) and Shu et al. (2025) but with an extra correction of $E(B-V)=0.045~\mathrm{mag}$ and $R_{{V}}=3.1$ corresponding to the Galactic extinction (Schlafly & Finkbeiner, 2011).
Figure 3: Radio emission of EP241021a and the fitting with a jet shock emission. The data are taken from Yadav et al. (2025), Gianfagna et al. (2025), and Shu et al. (2025), which are represented by squares, circles, and hexagrams, respectively. Despite sharing the same color labeling, data from different references may exhibit subtle frequency differences. Accordingly, theoretical curves are plotted over the corresponding frequency ranges, with model parameters sampled from their posterior distributions.
@@ 3.1 An explosion transient emission
Prior to detailed modeling of the multi-wavelength light curves of EP 241021a, we first conduct an order-of-magnitude analysis of the optical bump appearing at 10 days. First of all, the peak bolometric luminosity of the bump can be calculated to $L_{\rm p}\sim 10^{44}\mathrm{erg~s^{-1}}$ with a black body assumption which will be checked in Section 4. The blackbody temperature can be estimated to be within the range of $T=(1+z)h\nu/(2.82k_{\mathrm{B}})=(0.7\sim 1.7)\times 10^{4}~\mathrm{K}$, where the spectral peak frequency is considered to be located in the observational range of $(4\sim 10)\times 10^{14}~\mathrm{Hz}$, $k_{\mathrm{B}}$ and $h$ are the Boltzmann constant and the Planck constant, respectively. Then, the radius of the emitting photosphere can be estimated as
@@
[…]
@@
[…]
@@
where the initial value of the spin-down luminosity and the spin-down timescale are expressed as usual as $L_{\mathrm{sd,i}}=10^{47}B_{\mathrm{p,14}}^{2}P_{\mathrm{i,-3}}^{-4}~\mathrm{erg~s^{-1}}$ and $t_{\mathrm{sd}}=2\times 10^{5}~(1+z)B_{\mathrm{p,14}}^{-2}P_{\mathrm{i,-3}}^{2}~\mathrm{s}$ with $B_{\mathrm{p}}$ and $P_{\mathrm{i}}$ being the magnetic field strength and initial spin period of the magnetar, respectively. The convention $Q_{x}=Q/10^{x}$ is adopted in cgs units. In our calculations, we take $L_{\mathrm{sd,i}}$ and $t_{\mathrm{sd}}$ as free parameters directly. The coefficient $\xi=1-\exp(-\tau_{\mathrm{x}})$ is introduced to represent the fraction of the spin-down energy absorbed by the explosion ejecta, where $\tau_{\mathrm{x}}=3\kappa_{\mathrm{x}}M_{\mathrm{ej}}/(4\pi v_{\mathrm{ej}}^{2}t^{2})$ is the optical depth for X-rays with $\kappa_{\mathrm{x}}$ being the effective X-ray opacity. It is considered that the spin-down energy released from the magnetar is initially in the form of a Poynting flux and subsequently carried out by an ultra-relativistic electron-positron wind of a Lorentz factor $\sim 10^{3-7}$ (Kennel & Coroniti, 1984; Kirk et al., 2009; Bucciantini et al., 2011; Amato, 2020). This energy conversion occurs mainly through magnetic reconnection processes, which can simultaneously produce X-ray emission (Giannios & Spruit, 2005; Beniamini & Piran, 2014; Sironi & Spitkovsky, 2014; Kagan et al., 2015). When the magnetar wind catches up and collides with the previous ejecta, a reverse shock can be formed in the wind to convert its kinetic energy into heat. The expansion velocity of the reverse-shocked wind (i.e., the magnetar wind nebula; MWN) is continuous with that of the ejecta at the contact discontinuity. The hot electron-positron plasma in the MWN can release its internal energy through the synchrotron and inverse-Compton processes of relativistic leptons. The emitted photons would be mainly in X-rays too (Metzger et al., 2014; Yu et al., 2019), at least, at early times. Therefore, the X-ray opacity is involved in the absorption coefficient.
@@
$L_{\mathrm{th}}=e^{-(t/t_{\mathrm{diff}})^{2}}\times\int^{t}_{0}2L_{\mathrm{spl}}\left(t^{\prime}\right)\frac{t^{\prime}}{t^{2}_{\mathrm{diff}}}e^{(t^{\prime}/t_{\mathrm{tiff}})^{2}}dt,$ (4)
@@
Here, it should be pointed out that, as the ejecta expands, the X-ray emission generated from the MWN can increasingly penetrate the ejecta to be detected. Thus, a rebrightening of the X-ray emission accompanying the optical bump can be naturally understood. Such a feature makes EP241021a intrinsically different from the other analogous transient events such as EP240414a and EP250108a. In other words, the modeling of the X-ray emission of EP241021a would play a crucial role in identifying its origin. Additionally, if the total spin-down energy can finally exceed the initial kinetic energy of the ejecta, then it needs to be taken into account that the ejecta velocity $v_{\rm ej}$ can increase with time.